### v1

@@ 4.2 Implementation Details
Training. All models were trained on a single Nvidia A40 GPU with 42GB of VRAM for $100$ epochs, resulting in a training time of five days, with an Adam optimizer [34], a learning rate of $-55\mathrm{e}\text{-}$ 5, and a cosine scheduler [50] with 1000 warmup steps. One epoch consists of 1900 steps, each comprising a new scene and three randomly sampled views as supervision, scaled to $64\times 48$ pixels. Using more than three supervising views per scene leads to diminishing returns; using only three views thus reduces the total training time, a fact also observed by others [27]. Ray-based sampling allows our model trained on low-resolution images to render high-resolution outputs during inference. We sample 64 coarse followed by 64 fine points during training.
Inference. A forward pass to obtain the parameterized triplane is completed in 395 ms. Rendering an image of size $400\times 300$ from this triplane, using 128 uniformly-spaced points along each ray, takes 520 ms without the Projected Image Features (PIFs) vs 1461 ms when using PIFs. To improve the quality, a second pass based on the density of the first uniform sampling can be done by resampling 128 points in denser regions. This brings the rendering speed to 955ms without PIFs and 2853 ms with PIFs. Adding or removing the PIFs and sampling in a single or double fashion thus offers a trade-off between visual fidelity and speed.
@@ 4.3 Quantitative Results
We evaluate our method (6Img-to-3D) against several recent baselines such as NeRF [55], K-Planes [21] and the related few-shot method PixelNeRF [103] which Neo360 [30] also evaluates against.
Table 1: Performance comparison.
@@
TriPlane Overfitting 22.902 0.792 0.450
NeRF [55] 8.473 0.498 0.820
K-Planes [21] 9.307 0.437 0.773
PixelNeRF [104] 15.263 0.695 0.673
6Img-to-3D (Ours) 18.864 0.733 0.453
Though Neo360’s approach only targets inward-facing camera setups with largely overlapping cameras, it would be interesting to see how well Neo360 can handle our more difficult use-case of outward-facing cameras with minimal overlap and compare against it. However, it requires a considerable computational budget for training, i.e., 8 A100 GPUs for one day, which was not available in the work for this paper. Since this paper focuses on methods that can be trained efficiently on a single GPU and [30] does not allow straightforwardly to down their model, we could not compare against it. Furthermore, Neo360’s approach targets an inward-facing camera setup with widely overlapping camera views as input, whereas we only consider ego-vehicle views with outward-facing cameras. Their NERDS360 dataset [30] does not provide such a setup. Performance is measured using the established metrics PSNR, LPIPS [110], and MSE on the validation set. The results can be found in Tab. 1 and are visualized in Fig. 5.
We evaluated NeRF and K-Planes by training them on a subset of the evaluation set. The overfitted triplane was trained per evaluation scene on $80\%$ of the supervising spherical images to determine the obtainable upper bound when using the triplane representation. PixelNeRF was trained and evaluated under conditions equivalent to our method. Due to the negligible camera view overlap, NeRF and K-Planes cannot reconstruct the scene, emphasizing the task’s difficulty.
Figure 5: Visualization of the Comparison. PixelNeRF visibly struggles to reconstruct details and cannot resemble texture properly. Our method represents shape and appearance faithfully.
@@ 4.4 Ablation Study
Table 2: Ablation Results.
@@
w/o Scene Contraction 17.312 17.493 0.706 0.726 0.492 0.479
w/o $\mathcal{L}_{\text{LPIPS}}$ 18.835 18.953 0.723 0.736 0.555 0.538
w/o Proj. Image Features 18.396 18.440 0.719 0.726 0.484 0.488
w    SwinFIR $**{}^{\text{*}}$ 19.188 - 0.746 - 0.444
6Img-to-3D (Ours) 18.762 18.864 0.719 0.733 0.453 0.453
* Not included in the full model since it visually impairs image quality.
We show the effectiveness of our model elements in Tab. 2 and visualize our ablations in Fig. 6. All model variants are trained for 100 epochs and evaluated on the holdout validation set consisting of 100 scenes from town two. We find that the scene contraction combined with the contracted cross-attention mechanism substantially improves the model performance. The PIFs and LPIPS help adjust the color values further. While LPIPS leads to more human-appealing images, this comes at the cost of reduced PSNR performance. Double and single sampling here refers to rendering the image with or without a more fine-grained second sampling step.
Additionally, an off-the-shelf upscaler, SwinFIR [107], is trained and tested on the output of our 6Img-to-3D model. While slightly improving on all metrics, a visual analysis shows that the upscaler only outputs a smoothed image instead of adding otherwise missing details.
4.5 Qualitative Results
Depth Values. We visualize the expected termination depth of each ray, resulting in the depth maps shown in Fig. 7. While never trained on depth information explicitly, 6Img-to-3D outputs reasonable depth maps.
Figure 7: Depth Visualization. 6Img-to-3D infers a full 3D representation of the scene and can thus also be queried for density only, resulting in reasonable depth maps.
Consistency Over Timesteps. Our model works deterministically and thus obtains identical outputs for identical inputs. 6Img-to-3D also produces temporally consistent images across multiple consecutive timesteps without being explicitly trained to do so. A small change in the pose of the six input cameras results in a small change in the resulting output representation, visualized within the short trajectory in Fig. 8. When fed with six images per timestep, our model can also visualize longer scenes for example from a bird’s-eye view or a third-person perspective, see Fig. 8 long trajectory.
@@ 4.5 Qualitative Results
Figure 8: Trajectories. Consecutive timesteps are visualized with images from a BEV camera and behind and above the ego vehicle.
@@ 4.6 Effects of Scaling
Since our framework is self-supervised, we investigated whether our pipeline would benefit from a larger dataset. As shown in Fig. 9, the performance scales with the number of scenes.
@@
Figure 9: Effect of number of training scenes on performance.
5 Discussion
Conclusion. In this paper, we present 6Img-to-3D, an efficient and scalable method for reconstructing 3D representations of unbounded outdoor scenes from a few images. Our evaluation demonstrates that 6Img-to-3D can faithfully reconstruct third-person vehicle perspectives and birds-eye views using only six images with minimal overlap and without depth information. We achieve this by leveraging a triplanar representation of space, deformable attention mechanisms, feature projection and scene contraction. Our pipeline benefits from utilizing the features from the input images at multiple pipeline stages. A novel cross-attention sampling mechanism for contracted planes essentially drives performance gains. Our model could be further improved by additional training and using more data. We conclude this from the absence of overfitting during training and from how the model scales with data. An increased model size allows for a higher triplane resolution, which might help alleviate blurry texture for fine-grained details. Our model has around $\frac{1}{7}$ the amount of parameters of LRM [27], so visual fidelity could certainly be improved by scaling the model. We demonstrate the potential of our approach, particularly for embedded devices within vehicles, given its computational requirements. Seamless 2D-to-3D reconstruction could enhance driver assistance and autonomous driving systems, ultimately increasing road safety and navigation capabilities.
Future Directions. We see potential to improve the performance of our model. Initial experiments with an out-of-the-box upsampler trained on the data for post-rendering upscaling improved PSNR, SSIM, and LPIPS scores but introduced blurring (see Fig. 6). Incorporating the six source image information into the upscaling process represents a potentially valuable research direction to improve upscaling performance. The rendered could be extended to incorporate the viewing direction if incorporating view-dependent effects is required. This would allow the model to accurately recreate real-world materials, such as shiny metals or wet surfaces, with specular or glossy properties. We made some first attempts to run our trained model zero-shot on the nuScenes dataset [9]. Those zero-shot tests showed promising results that the model can be extended to real-world data. We assume that real-world driving data could be utilized by leveraging multi-timestep images for training. Using LiDAR for depth supervision or implementing a separate segmentation head would be natural extensions of our work.

### v2

@@ 4.2 Implementation Details
Training. All models were trained on a single Nvidia A40 GPU with 42GB of VRAM for $100$ epochs, resulting in a training time of five days, with an Adam optimizer [40], a learning rate of $-55\mathrm{e}\text{-}$ 5, and a cosine scheduler [55] with 1000 warmup steps. One epoch consists of 1900 steps, each comprising a new scene and three randomly sampled views as supervision, scaled to $64\times 48$ pixels. Using more than three supervising views per scene leads to diminishing returns; using only three views thus reduces the total training time, a phenomenon also observed by others [29]. Ray-based sampling allows our model trained on low-resolution images to render high-resolution outputs during inference. We sample 64 coarse followed by 64 fine points during training.
Inference. A forward pass to obtain the parameterized triplane is completed in 395 ms. Rendering an image of size $400\times 300$ from this triplane, using 128 uniformly-spaced points along each ray, takes 520 ms without the Projected Image Features (PIFs) vs 1461 ms when using PIFs. To improve the quality, a second pass based on the density of the first uniform sampling can be done by resampling 128 points in denser regions. This brings the rendering speed to 955ms without PIFs and 2853 ms with PIFs. Adding or removing the PIFs and sampling in a single or double fashion thus offers a trade-off between visual fidelity and speed. While being trained on a 40GB GPU, our method requires as little as 10.5 GB of vRAM during inference.
@@ 4.3 Quantitative Results
Figure 5: Qualitative Results. PixelNeRF and SplatterImage visibly struggle to model details and cannot resemble texture properly. Our method represents shape and appearance faithfully. 6Img-to-3D infers a full 3D representation of the scene and can thus also be queried for density only, resulting in reasonable depth maps. Due to their low performance, we do not visualize SplatFacto at the top and SplatFacto K-Planes and NeRFacto at the bottom part.
We evaluate our method (6Img-to-3D) against several approaches such as SplatFacto [103], K-Planes [23], NeRFacto [81], the related few-view methods PixelNeRF [105] and the recent SplatterImage [77]. We also compare against two metric monocular depth estimation techniques, namely ZoeDepth [7] and Metric3D [30].
Table 1: Result Comparison. Due to occlusion artifacts we also compute all metrics for ZoeDepth and Metric3D while masking out regions occluded without points, indicated with a ‡. For a fair comparison, the unmasked values are considered when ranking the methods. Colors indicate first, second, and third best values.
@@
ZoeDepth [7] 5.466 0.254 0.563 11.728
${}^{\text{\textdaggerdbl}}$ [7] 14.202 0.661 0.292 9.378
Metric3D [30] 6.314 0.296 0.554 10.049
${}^{\text{\textdaggerdbl}}$ [30] 13.699 0.600 0.336 8.655
NeRFacto [81] 10.943 0.298 0.791
K-Planes [23] 11.356 0.463 0.633
SplatFacto [103] 11.607 0.486 0.658
PixelNeRF [106] 14.500 0.550 0.652 19.235
SplatterImage [77] 17.791 0.580 0.568 11.049
6Img-to-3D 18.683 0.726 0.451 6.232
We would have liked to compare against Neo360’s approach, however, the method requires a considerable computational budget for training, i.e., 8 A100 GPUs, which was not available for this paper. Performance is measured using the established metrics PSNR, LPIPS [112], and MSE on the test set. The results can be found in Tab. 1. We evaluated SplatFacto, K-Planes, and NeRFacto by training them on a subset of the evaluation set. Due to the negligible camera view overlap, the iterative methods cannot reconstruct the scene. PixelNeRF and SplatterImage are trained and evaluated under conditions equivalent to our method. To evaluate Metric3D and ZoeDepth, monocular depth maps are generated from ego images and then unprojected into space to create a colored point cloud, from which novel views can be rendered. For details of the evaluation protocols, see the appendix.
4.4 Qualitative Results
Novel View Visualization. When comparing the visual results we find our method to be visually most appealing. Novel views resulting from projecting monocular depth values into space are often misaligned and unobserved parts lead to visible artifacts. PixelNerf and SplatterImage struggle to model texture and geometry faithfully. Since our method performs single-shot inference over occluded or unobserved parts of the scene those inferred parts are less sharp than in ground truth images. Results are visualized in the upper part of Fig. 5.
Depth Visualization. We visualize the expected termination depth of each ray, resulting in the depth maps shown in the bottom part of Fig. 5. For the colored point clouds resulting from the unprojected depth images, the closest point along the viewing direction per pixel is rendered. While never trained on depth information explicitly, 6Img-to-3D outputs reasonable depth maps for novel views.
@@ 4.5 Ablation Study
We show the effectiveness of our model elements in Tab. 2 and visualize our ablations in Fig. 6. All model variants are trained for 100 epochs and evaluated on the holdout validation set consisting of 100 scenes from town two. We find that the scene contraction combined with the contracted cross-attention mechanism substantially improves the model performance. The PIFs and LPIPS help adjust the color values further. While LPIPS leads to more human-appealing images, this comes at the cost of reduced PSNR performance. Double and single sampling here refers to rendering the image with or without a more fine-grained second sampling step. Additionally, we performed ablation studies on how well our architecture scales with data, plotted in appendix Sec. 6.6.
Table 2: Ablation Study. The colors indicate best, second-best, and third-best values.
@@
w/o SC 17.155 17.352 0.699 0.721 0.487 0.474
w/o $\mathcal{L}_{\text{LPIPS}}$ 18.674 18.812 0.718 0.731 0.551 0.537
w/o $f_{PIF}$ 18.201 18.256 0.714 0.718 0.482 0.487
6Img-to-3D 18.567 18.683 0.712 0.726 0.451 0.451
5 Discussion
Conclusion. In this paper, we present 6Img-to-3D, a novel, high-fidelity method for reconstructing 3D representations of unbounded outdoor scenes from a few images. Our evaluation demonstrates that 6Img-to-3D can faithfully reconstruct third-person vehicle perspectives and birds-eye views using only six images with minimal overlap and without depth information. We achieve this by leveraging a triplanar representation of space, deformable attention mechanisms, feature projection, and scene contraction. Our pipeline benefits from utilizing the features from the input images at multiple pipeline stages. A custom cross-attention sampling mechanism for contracted planes essentially drives performance gains. Unlike for example SplatterImage, our method is natively designed to incorporate multiple images at once and integrate their information into a single scene representation. We demonstrate the potential of our approach, as a single-shot high-fidelity novel view synthesis method for unbounded scenes. We consider this problem to be important for applications including the teleoperation of large vehicles (such as trucks) and vision-based parking assistance systems.
Future Directions. Since we are using querying-based color and density computation our rendering is less fast than for example Gaussian Splatting. Future work includes porting our rendering pipeline towards Gaussian Splats.
We made some first attempts to run our trained model zero-shot on the nuScenes dataset [10]. Those tests show promising results that the model can zero-shot be extended to real-world data, for detail see appendix Sec. 6.7. Using sim-to-real transfer methods could aid this process. We assume that real-world driving data could be utilized by leveraging multi-timestep images for training or fine-tuning whereby LiDAR could provide an additional depth supervision signal.
Further, tuning scene contraction and other hyperparameters used within our method could help reduce blurring artifacts. Similarly, additional data and longer training could improve performance. We conclude this from the absence of overfitting during training and from how the model scales with data. An increased model size allows for a higher triplane resolution, which might help alleviate blurry texture for fine-grained details. Our model has around $\frac{1}{7}$ the amount of parameters of LRM [29], so visual fidelity could certainly be improved by scaling the model.
During initial experiments with an upsampler trained on the data for post-rendering upscaling improved PSNR, and SSIM scores but introduced blurring. See appendix Sec. 6.10 for further information. Incorporating the six source image information into the upscaling process represents a potential research direction to improve performance.
Supplementary Material
6 Appendix
6.1 Bounded Scene Paramterization
The transformation from world coordinates $\bm{x_{W}}$ to grid coordinates $\bm{x_{grid}}$ can also be done when a bounded scene is given.
Bounded scenes: The world coordinates $\bm{x_{W}}$ are transformed into grid coordinates $\bm{x_{grid}}$ by scaling the scene with the scaling parameter $\bm{s}=[s_{h},s_{w},s_{z}]$ and applying the offset $\bm{o}=[o_{,}o_{w},o_{z}]$ according to Eq. 7 for element-wise operations. When depth maps are available, out-of-bound pixels of the training images can be masked with a white background during an initial testing phase, so the network learns only to reconstruct the inside of the bounded scene.
$\bm{x_{grid}}=\frac{\bm{x_{W}}-\bm{o}}{\bm{s}}$ (7)
6.2 Shape and Parameter Details
In our model description, we follow the PyTorch convention of Batch $\times$ Channel $\times$ Height $\times$ Width ($B\times C\times H\times W$). The shape of 6Img-to-3D is shown in Tab. 3. While the size of the input images during training remains fixed, the shape of the rendered images can be varied during both training and inference.
Table 3: Shape and parameter values of 6Img-to-3D trained on the introduced dataset.
Part Shape Parameters
Input Images $6\times 3\times 928\times 1600$ n.a.
ResNet Maps $6\times 512\times 116\times 200$ 44M
$6\times 1024\times 58\times 100$
$6\times 2048\times 29\times 50$
Feature Maps $6\times 128\times 116\times 200$ 1M
$6\times 128\times 58\times 100$
$6\times 128\times 29\times 50$
$6\times 128\times 15\times 25$
Attention Mechanism n.a. 16M
Triplane $128\times 200\times 200$ n.a.
$128\times 200\times 16$
$128\times 16\times 200$
MLP decoder input: 384 84K
output: 4
Output Image $3\times 600\times 800$
6.3 Baselines and Competing Methods
6.3.1 Depth unprojection
Figure 7: Depth Unprojection. A monocular depth method is used to obtain depth maps for the ego views (step A). Depth maps and RGB ego views are then used to obtain a colored point cloud (step B). From the colored point cloud novel views can be rendered (step C).
ZoeDepth [7] is a zero-shot metric depth estimation technique. The method is trained on multiple datasets, including the autonomous driving dataset KITTI. The method builds on the MiDaS depth estimation framework. We utilize the model’s publicly available code along with pre-trained weights that have been trained on both the NYU Depth v2 and KITTI datasets. Metric3D [30] is a metric 3D reconstruction method using a canonical camera space transformation method. The method can perform zero-shot metric depth and surface normal estimation from a single image. The publicly available code is used, and the largest version of the pre-trained model, vit_giant, is used.
Unprojection steps: Since our method reconstructs a Lambertian scene, we compare it against a depth baseline. We use one of the monocular depth estimation methods to obtain a depth map for each of the six ego input images resized to $842\times 842$ to fit the model. Since camera intrinsics and extrinsic are known, we can use the depth maps to project the image pixels into space to obtain a colored point cloud (sometimes also referred to as 2.5D). The obtained colored point cloud can now be used to rasterize novel exo views. The pipeline is visualized in Fig. 7. We optimized the available parameters (e.g., the size of the rendered points) to obtain the best possible results. We use open3d and pytorch3d to implement the pipeline.
6.3.2 SplatterImage
SplatterImage [77] is designed for inferring 3D Gaussian Splatting primitives from conditioning images in a pixel-aligned fashion. A U-net style image-to-primitive mapping network supported by a cross-attention mechanism maps the input RGB images to a ’Splatter Image’ containing opacity, position, shape, and color information of a 3D Gaussian. We use the repository released by the authors.
6.3.3 PixelNeRF
PixelNeRF [106] is a sparse novel view synthesis method. PixelNeRF weakens some of the shortcomings of the original NeRF paper by leveraging projected image features and training across multiple scenes. We use the re-implementation introduced in the code of Neo360 [33].
6.3.4 Iterative training methods
K-Planes [23] factorizes a 3D scene into multiscale planes. The plane features are learned using differentiable volume rendering, sampled using multiscale bilinear interpolation, and rendered using a small MLP. We use the hybrid version of the model and the GitHub users Giodiro’s reimplementation of the model available within NeRFstudio.
NeRFacto [81] is a combination of several published methods. The method is optimized to work particularly well for real data captures. The following techniques are combined in this method: camera pose refinement, per-image appearance embedding learning, proposal sampling, scene contraction, and hash encoding. We use the model as part of NeRFStudio.
SplatFacto [37] is a re-implementation of the original 3D Gaussian Splatting paper [37] within NeRFStudio. The method explicitly stores a collection of 3D volumetric Gaussians to parameterize the scene. During rendering the 3D Gaussions are ’splatted’ to obtain per-pixel colors. We use the model as part of NeRFStudio.
6.4 Training Details
K-Planes. We train each of the models for 30k steps on a single Tesla T4 GPU with 16GB of VRAM. We follow the model’s default NeRFStudio [81] settings for training. Near and far bounds of the scene are adjusted to 0.1 to 60 to best accommodate the scenes. Additionally, scene contraction is applied. The training took around 1.5 hours per model.
NeRFacto. We train each of the models for 30k steps on a single Tesla T4 GPU with 16GB of VRAM. We follow the model’s default NeRFStudio [81] settings for training. We disable the model’s use of an appearance embedding since those lead to problems during the evaluation, and we also deactivate the camera pose optimization because we already provide the model with ground truth poses. The near and far bounds are set to 0.1 and 60. Each model is trained for a total of 1 hour.
SplatFacto. We train each of the models for 30k steps on a single Tesla T4 GPU with 16GB of VRAM. We again follow the model’s default NeRFStudio [81] settings for training. The model took a total of 20 minutes to train.
PixelNeRF. We train PixelNeRF for 100k steps on a Nvidia A40 GPU with 42GB of VRAM, with an Adam optimizer [40] and a learning rate of $-31\mathrm{e}\text{-}$ 3. Total training time accumulates to five days.
SplatterImage. We train SplatterImage for five days across five 3090 GPUs with 24GB of VRAM. During training, the supervision images are scaled to $128\times 128$ pixels. We use the multi-input image variant of the model to accommodate all six input views.
ZoeDepth and Metric3D are not fine-tuned.
@@ 6.4 Training Details
Figure 8: Trajectory Visualization. Consecutive timesteps are visualized with images from a BEV camera and a view location behind and above the ego vehicle.
6.5 Consistency Over Timesteps
Our model works deterministically and thus obtains identical outputs for identical inputs. 6Img-to-3D also produces temporally consistent images across multiple consecutive timesteps without being explicitly trained to do so. A small change in the pose of the six input cameras results in a small change in the resulting output representation, visualized within the short trajectory in the upper part of Fig. 8. When fed with six images per timestep, our model can also visualize longer scenes for example from a bird’s-eye view or a third-person perspective, see the bottom part of Fig. 8 long trajectory.
@@ 6.6 Effects of Scaling
Since our framework is self-supervised, we investigated whether our pipeline would benefit from a larger dataset. As shown in Fig. 9, the performance scales with the number of scenes. When tested with 100, 200, 500, 1000, and 1500 scenes, we find that a higher number of scenes leads to improved model performance.
@@
Figure 9: Effect of number of training scenes on performance. Only the number of training scenes was varied, while all other parameters were kept constant.
6.7 Zero-shot nuScenes Performance visualization
We attempt an early analysis of our 6Img-to-3D applied zero-shot to nuScenes. The results are vizualized in Fig. 11. Without additional training, the model can display the appearance of nuScenes. Unfortunately, the camera setup within the SEED4D dataset dpes not match the nuScenes data exactly. Camera positions, FOV, and scene style (sunny vs. cloudy weather) differ, making zero-shot transfer difficult. The promising results motivate to extent the model to real-world data. Consistency across multiple timesteps is shown in Fig. 10.
Figure 10: nuScenes Short Trajectory. Consecutive timesteps inferred during zero-shot inference on nuScenes.
Figure 11: 6Img-to-3D with nuScenes Data. By inputting six images from nuScenes into our trained model pipeline, we can generate novel views.
6.8 Failure Cases
In Fig. 12, failure cases of our method are visualized. Our model struggles with zero-shot reconstructing fine details, as seen in examples one and two. This might be because we only trained the model with an image resolution of $64\times 48$ pixels. Partially, model failures can be attributed to the dataset itself. Some of the supervising images in the training set are located inside buildings. Within the Carla simulator, walls are sometimes transparent from within the building, which leads the model to pick up on this physically impossible artifact. Those failure cases are shown in examples three and four. Example five shows an instance where the model is unable to infer the structure of the building beyond what is contained in the input images.
Figure 12: Visualization of Failure Cases. Failure cases can partially be attributed to the CARLA engine, which sometimes struggles to render walls when within a building.
6.9 Extended Input-Output Visualization
To further illustrate the task our model is handling, we visualize several additional input images and example novel views in Fig. 13. Ego-centric input views are characterized by minimal view overlap, an unbounded line of sight, and frequent occlusions. Both the texture and geometry of objects in a scene vary widely. The exo-centric views represent a diverse mixture of non-ego views. The ego vehicle is not included in the ground truth novel views and, therefore, does also not appear in the rendered novel views.
Figure 13: Input-Output Examples. During inference time, 6Img-to-3D takes six surround vehicle RGB images with a small overlap as input and returns a parameterized triplane from which arbitrary novel views can be rendered.
6.10 Extended Qualitative Results Comparison
We extend the visualization from our ablations in Fig. 14 to Fig. 16 to better understand our model’s performance. The results are not cherry-picked. Additionally, an off-the-shelf upscaler, SwinFIR [109], is trained and tested on the output of our 6Img-to-3D model. While slightly improving on all metrics (PSNR 19.188 up from 18.683, SSIM increased to 0.746 from 0.726, and LPIPS decreased to 0.444 from 0.451), a visual analysis shows that the upscaler only outputs a smoothed image instead of adding otherwise missing details. We thus decided not to include the upscaled results in the main analysis.
Figure 14: Qualitative Ablation Results (I).
Figure 15: Qualitative Ablation Results (II).
Figure 16: Qualitative Ablation Results (III).
6.11 Extended Qualitative Results
This section provides additional qualitative results of reconstructed views and their associated depth maps for various methods. The depictions can be found in Fig. 17 for the rendered images and in Fig. 18 for the depth maps.
Figure 17: Extended Qualitative Results.
Figure 18: Extended Qualitative Results. Due to their low performance, we do not visualize SplatFacto at the top and SplatFacto K-Planes and NeRFacto at the bottom part.
6.12 Dataset Visualization
Towns differ in buildings (residential and commercial), types of bridges, the number of rivers, the number of traffic lights, and the kind of junctions. All towns contain pedestrians. For further details about the dataset, please see the dataset publication [36] or the additional visualization we provide in Sec. 6.12.
Figure 19: Samples from the SEED4D Dataset. Ego- and Exo-centric images for a number of towns.